# Jacquet modules of Tate cohomology and base change lifting

### v1

@@ 3.1. Twisted Jacquet modules
Let $(\pi,V)$ be a smooth $\Lambda[G(F)]$ module, where $\Lambda$ is the ring of integers of the maximal unramified extension of $\mathbb{Q}_{\ell}$ in $\overline{\mathbb{Q}}_{\ell}$. Let $P=MN$ be an $F$-parabolic subgroup of $G$, where $M$ is a Levi subgroup of $P$ and $N$ is the unipotent radical of $P$. Let $\Theta:N(F)\rightarrow\Lambda^{\times}$ be a character. Let $V(N(F),\Theta)$ be the $\Lambda$-module generated by the following set of vectors
@@
The quotient space $V/V(N(F),\Theta)$ is called a twisted Jacquet module of $V$, which is denoted by $V_{N(F),\Theta}$. This space is characterized by the set of elements $v\in V$ for which there exists some compact open subgroup $\mathcal{N}_{v}$ in $N(F)$ such that
@@
where $dn$ is a Haar measure on $N(F)$. The space $V_{N(F),\Theta}$ is a smooth representation of the $M(F)$ stabiliser of the character $\psi$, to be denoted by $M(F)_{\Theta}$. In particular, when $\Theta=1$, the trivial character of $N(F)$, then the smooth $\Lambda[M(F)]$ module $V_{N(F),1}$ is called the Jacquet module of $V$. We also use the notation $r_{N(F)}(V)$ for $V_{N(F),1}$.
@@ 3.2. Integral representation and $-\ell$ reduction
Let $\sigma$ be an automorphism of $G$ of order $\ell$, where $\ell$ is a prime different from the residue characteristic of $F$. Let $P$ be a $\sigma$ stable $F$-parabolic subgroup of $G$, and let $N$ be the unipotent radical of $P$. Note that $N$ is $\sigma$-stable and let $M$ be a $\sigma$-stable Levi subgroup $P$ (for the existence of $M$, see [KP23, Lemma 12.5.4] for a reference). Given any parabolic subgroup $Q$ of $G^{\sigma}$, there exists a $\sigma$-stable parabolic subgroup $P$ of $G$ such that $Q=P^{\sigma}$ (see [KP23, Lemma 12.5.4] for a reference). Note that $U=N^{\sigma}$ is the unipotent radical of $Q$ and $M^{\sigma}$ is a Levi subgroup of $Q$. Let $\Theta:N(F)\rightarrow\Lambda^{\times}$ be a character such that $\Theta^{\sigma}=\Theta$.
Lemma 3.1.
Let $(\pi,\mathcal{L})$ be a smooth $\Lambda[G(F)\rtimes\langle\sigma\rangle]$ module. For each $i\in\{0,1\}$, the image of the natural map $\widehat{H}^{i}(\mathcal{L}(N(F),\Theta))\rightarrow\widehat{H}^{i}(\mathcal{L})$ is equal to $\widehat{H}^{i}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$, where $\overline{\Theta}$ is the the $-\ell$ reduction of $\Theta$.
@@
We first assume that $i=0$. Let $v$ be in the image of $\varphi_{0}$, and let $\tilde{v}$ be a lift of $v$ in $\mathcal{L}(N(F),\Theta)^{\sigma}$. There exists a compact open subgroup $\mathcal{N}\subseteq N(F)$ such that
@@
[…]
@@
Since the action of $\langle\sigma\rangle$ on $X\setminus X^{\sigma}$ is free, there exists a subset $\mathcal{U}\subseteq X\setminus X^{\sigma}$ such that $X\setminus X^{\sigma}$ is the disjoint union of $\sigma^{i}\mathcal{U}$ for $1\leq i\leq\ell$. As $\tilde{v}$ is a $\langle\sigma\rangle$-fixed vector, we get
@@
where ${\rm Nr}=1+\sigma+\cdots+\sigma^{\ell-1}$. This shows that
@@
This implies that $v$ belongs to the space $\widehat{H}^{0}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Conversely, let $w$ be an element of $\widehat{H}^{0}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Let $\mathcal{N}_{\sigma}$ be the compact open subgroup of $N^{\sigma}(F)$ with
@@
Now, consider the case $i=1$. Let $v$ be an element in the image of $\varphi_{1}$. Let $\tilde{v}$ be a lift of $v$ in ${\rm Ker}({\rm Nr}_{|\mathcal{L}(N(F),\Theta)})$. Then there exists $\langle\sigma\rangle$ stable compact open subgroups $\mathcal{N}^{\prime}\subset\mathcal{N}$ such that $\mathcal{N}^{\prime}$ fixes the vector $\tilde{v}$, $\Theta$ is trivial on $\mathcal{N}^{\prime}$, and
@@
[…]
@@
Thus, we get that $\sum_{x\in X^{\sigma}}\overline{\Theta}^{-1}(x)\,\pi(x)v$ is trivial in $\widehat{H}^{1}(\mathcal{L})$. Conversely, let $v$ be an element in $\widehat{H}^{1}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Let $\tilde{v}\in{\rm Ker}({\rm Nr})$ be a lift of $v$. Consider the element
@@
Theorem 3.2.
Let $P$ be a $\sigma$ stable parabolic subgroup of $G$, and let $N$ be the unipotent radical of $P$. Let $\mathcal{L}$ be a smooth $\Lambda[G(F)\rtimes\langle\sigma\rangle]$ module. The natural map $\widehat{H}^{i}(\mathcal{L})_{N^{\sigma}(F),\overline{\Theta}}\rightarrow\widehat{H}^{i}(\mathcal{L}_{N(F),\Theta})$ is injective.
@@
The long exact sequence of Tate cohomology corresponding to the above short exact sequence is equal to:
@@
Using Lemma 3.1, the image of $\varphi_{i}$ is equal to $\widehat{H}^{i}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Hence, we get the theorem. ∎
@@
Corollary 3.3.
Let $G$ be a connected reductive algebraic group defined over $F$, and let $\sigma$ be an automorphism of $G$ of order $\ell$, where $\ell$ and $p$ are distinct primes. Let $\Pi$ be a $\langle\sigma\rangle$ equivariant cuspidal $\ell$-modular representation of $G(F)$. Then each Tate cohomology $\widehat{H}^{i}(\Pi)$ is a cuspidal representation of $G^{\sigma}(F)$.
@@
Let $Q$ be a proper $F$-parabolic subgroup of $G^{\sigma}$. By [KP23, Lemma 12.5.4], there exists a proper $\sigma$ stable $F$-parabolic subgroup $P$ of $G$ such that $P^{\sigma}=Q$. Let $N$ be the unipotent radical of $P$. Then $N^{\sigma}$ is equal to the unipotent radical of $Q$, to be denoted by $U$. Now, using Theorem 3.2 for $\Theta=1$, the trivial character of $N(F)$, we get an injection
@@
In the paper [TV16, Conjecture 6.1], the authors conjectured that if $\Pi$ is an irreducible $\ell$-modular representation of $G(F)$ that is $\langle\sigma\rangle$ equivariant, which means $\Pi$ extends to a representation of $G(F)\rtimes\langle\sigma\rangle$, then each Tate cohomology group $\widehat{H}^{i}(\Pi)$, $i\in\{0,1\}$, has finite length as a representation of $G^{\sigma}(F)$. In our previous work [ND24, Proposition 5.5], this conjecture is proved for $G={\rm GL}_{n}$ in the context of local base change. In this subsection, we generalise this result for arbitrary connected reductive groups.
@@
Let $E/F$ be a finite Galois extension with $[E:F]=\ell$, where $\ell$ and $p$ are distinct primes. We fix a generator $\sigma$ of the cyclic group ${\rm Gal}(E/F)$. Let $G$ be a connected reductive group defined over $F$. Fix a minimal $F$-parabolic subgroup $P_{0}$ of $G$. Let $\Omega$ denote the set of all proper $F$-parabolic subgroups of $G$ containing $P_{0}$. For $P\in\Omega$, we denote by $U$ the unipotent radical of $P$. Let $\pi$ be a smooth $R$-representation of $G(F)$, where the coefficient field $R$ is either $\overline{\mathbb{Q}}_{\ell}$ or $\overline{\mathbb{F}}_{\ell}$. We have the natural map obtained from Frobenius reciprocity:
$\pi\longrightarrow\prod_{P\in\Omega}i_{P(F)}^{G(F)}\circ(\pi_{U(F)}),$ (3.3)
@@
$\pi_{c}\longrightarrow\prod_{P\in\Omega}i_{P(F)}^{G(F)}((\pi_{c})_{U(F)}).$ (3.4)
Since the functors $i_{P(F)}^{G(F)}$ preserves injective maps and $r_{U(F)}$ is an exact functor, we get a contradiction to the fact that $\pi_{c}$ is the kernel of the map (3.3). Thus, $\pi_{c}$ is a cuspidal representation. Using this observation, we now prove the following finiteness result.
Theorem 3.4.
Let $E/F$ be a finite Galois extension of prime degree $\ell$, where $\ell\neq p$. Let $G$ be a connected reductive group defined over $F$. Let $\Pi$ be a finite length ${\rm Gal}(E/F)$ equivariant $\ell$-modular representation of $G(E)$. Then, the Tate cohomology groups $\widehat{H}^{i}({\rm Gal}(E/F),\Pi)$ are finite length $\ell$-modular representations of $G(F)$.
@@
We set $\Gamma={\rm Gal}(E/F)$ and $\sigma$ is the Frobenius element. We use induction on the length of $\Pi$ and the dimension of $G$ to prove the theorem. All Tate cohomology groups are with respect to the action of ${\rm Gal}(E/F)$ and we drop it from the notation of Tate cohomology. Suppose $\Pi$ is irreducible. Consider the natural map
$0\longrightarrow\widehat{H}^{i}(\Pi)_{c}\longrightarrow\widehat{H}^{i}(\Pi)\longrightarrow\prod_{P\in\Omega}i_{P(F)}^{G(F)}(\widehat{H}^{i}(\Pi)_{U(F)})$ (3.5)
For each parabolic $P\in\Omega$ with unipotent radical $U$ and a Levi subgroup $M$, the natural map gives an $M(F)$ equivariant injective map (Theorem 3.2)
$\widehat{H}^{i}(\Pi)_{U(F)}\hookrightarrow\widehat{H}^{i}(\Pi_{U(E)}).$
Since $\Pi_{U(E)}$ is a finite length representation of $M(E)$ and ${\rm dim}(M)<{\rm dim}(G)$, the Tate cohomology group $\widehat{H}^{i}(\Pi_{U(E)})$ is a finite length representation of $M(F)$ (by induction hypothesis). This implies that $\widehat{H}^{i}(\Pi)_{U(F)}$ is a finite length $M(F)$ representation, and so is $i_{P(F)}^{G(F)}(\widehat{H}^{i}(\Pi)_{U(F)})$ as the parabolic induction preserves finite length ([Vig96, Chapter II, Section 5.13]). Thus, the representation
$\prod_{P\in\Omega}i_{P(F)}^{G(F)}(\widehat{H}^{i}(\Pi)_{U(F)})$
is of finite length. Now, note that the center $Z(E)$ of $G(E)$ acts on the space of $\Pi$ by a character (as $\Pi$ is irreducible). We deduce that the center $Z(F)$ of $G(F)$ acts on $\widehat{H}^{i}(\Pi)$ by a character. Hence, the cuspidal $G(F)$-representation $\widehat{H}^{i}(\Pi)_{c}$, being a subrepresentation of $\widehat{H}^{i}(\Pi)$, admits a central character. Since $\widehat{H}^{i}(\Pi)$ is finitely generated as a representation of $G(F)$ (Theorem 2.3), it follows from [Dat09, Corollary 4.5]) that $\widehat{H}^{i}(\Pi)_{c}$ is also finitely generated. Finitely generated $\ell$-modular cuspidal representation of $G(F)$ with central character is of finite length, the Tate cohomology $\widehat{H}^{i}(\Pi)_{c}$ is a finite length representation. We conclude that the $\ell$-modular $G(F)$-representation $\widehat{H}^{i}(\Pi)$ is of finite length.
Now, we consider the general case where $\Pi$ is not necessarily an irreducible representation of $\overline{\mathbb{F}}_{\ell}[G(F)]$. Let $\Pi_{1}$ be an irreducible sub-representation of $\Pi$ and set $\Pi_{1}^{\prime}$ to be the sub-representation of $\Pi$ spanned by $\Pi_{1}^{\sigma^{i}}$, for $0\leq i<\ell$. Note that $\Pi_{1}^{\prime}$ is $\sigma$-stable, and we consider the short exact sequence of $\overline{\mathbb{F}}_{l}[G(E)\rtimes\langle\sigma\rangle]$ modules
@@
3.5. Genericity of Tate cohomology
For the rest of this subsection, $G$ will be a quasi-split connected reductive group defined over $F$. We follow the notations defined in Subsection 3.4.1.
3.5.1.
Let $B$ be a Borel subgroup of $G$ defined over $F$. Let $S$ be a maximal $F$-split torus and let $\Phi$ the set of roots of $S$ in $G$, and let $\Delta$ be the set of simple roots in $\Phi$ for the choice of $B$. Let $\mathfrak{g}$ be the Lie algbera of $G$ and let $\mathfrak{g}_{\alpha}$ be the subspace of $\mathfrak{g}$ for which $S$ acts via $\alpha$. The root subgroup corresponding to $\alpha\in\Delta$ is denoted by $U_{(\alpha)}$, this the unique connected subgroup of $G$ whose Lie algebra is $\mathfrak{g}_{\alpha}\oplus\mathfrak{g}_{2\alpha}$ ($\mathfrak{g}_{2\alpha}$ is taken to be zero when $2\alpha\not\in\Phi$). Let $R$ denote $\overline{\mathbb{Q}}_{\ell}$ or $\overline{\mathbb{F}}_{\ell}$. A character $\Theta:U(F)\rightarrow R^{\times}$ is called non-degenerate whenever $\Theta$ is non-trivial on $U_{(\alpha)}(F)$. Note that
$U(E)/U(E)_{\rm der}\simeq\prod_{\alpha\in\Delta}U_{(\alpha)}(E)/U_{(2\alpha)}(E).$
Here, $E$ is a Galois extension of $F$ with $[E:F]=\ell$. Note tha the quotient $U_{(\alpha)}/U_{(2\alpha)}$ is $F$ is isomorphic to an $F$-vector space denoted by $V_{\alpha}$. Thus we choose a non-degenerate character $\Theta_{E}:U(E)\rightarrow R^{\times}$ which is ${\rm Gal}(E/F)$ stable and the restriction of $\Theta_{E}$ to $U(F)$ is denoted by $\Theta_{F}$ (which is a non-degenerate character of $U(F)$). We set

### v2

@@ 3.1. Twisted Jacquet modules
Let $(\pi,V)$ be a smooth $\Lambda[G(F)]$ module, where $\Lambda$ is the ring of integers of the maximal unramified extension of $\mathbb{Q}_{\ell}$ in $\overline{\mathbb{Q}}_{\ell}$. Let $P=MN$ be an $F$-rational parabolic subgroup of $G$, where $M$ is an $F$-rational Levi subgroup of $P$ and $N$ is the unipotent radical of $P$. Let $\Theta:N(F)\rightarrow\Lambda^{\ast}$ be a character. Let $V(N(F),\Theta)$ be the $\Lambda$-module generated by the following set of vectors
@@
The quotient space $V/V(N(F),\Theta)$ is called a twisted Jacquet module of $V$, which is denoted by $V_{N(F),\Theta}$. This space is characterised by the set of elements $v\in V$ for which there exists some compact open subgroup $\mathcal{N}_{v}$ of $N(F)$ such that
@@
where $dn$ is a Haar measure on $N(F)$. The space $V_{N(F),\Theta}$ is a smooth representation of the $M(F)$ stabiliser of the character $\Theta$, to be denoted by $M(F)_{\Theta}$. In particular, when $\Theta=1$, the trivial character of $N(F)$, then the smooth $\Lambda[M(F)]$ module $V_{N(F),1}$ is called the Jacquet module of $V$. We also use the notation $r_{N(F)}(V)$ for $V_{N(F),1}$.
@@ 3.2. Integral representation and $-\ell$ reduction
Let $\sigma\in{\rm Aut}(G)$, defined over $F$, be an element of order $\ell$, where $\ell$ is a prime different from $p$. Let $\Gamma$ be the group generated by $\sigma$. Let $P$ be a $\sigma$-stable $F$-parabolic subgroup of $G$, and let $N$ be the unipotent radical of $P$. Note that $N$ is $\sigma$-stable and let $M$ be a $\sigma$-stable Levi subgroup $P$ (for the existence of $M$, see [KP23, Lemma 12.5.4] for a reference). Given any $F$-parabolic subgroup $Q$ of $G^{\sigma}$, there exists a $\sigma$-stable $F$-parabolic subgroup $P$ of $G$ such that $Q=P^{\sigma}$ (see [KP23, Lemma 12.5.4] for a reference). Note that $U=N^{\sigma}$ is the unipotent radical of $Q$ and $M^{\sigma}$ is a Levi subgroup of $Q$. Let $\Theta:N(F)\rightarrow\Lambda^{\ast}$ be a character such that $\Theta^{\sigma}=\Theta$.
Lemma 3.3.1.
Let $(\pi,\mathcal{L})$ be a smooth $\Lambda[G(F)\rtimes\Gamma]$ module. For each $i\in\{0,1\}$, the image of the natural map $\widehat{H}^{i}(\mathcal{L}(N(F),\Theta))\rightarrow\widehat{H}^{i}(\mathcal{L})$ is equal to $\widehat{H}^{i}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$, where $\overline{\Theta}$ is the the $-\ell$ reduction of $\Theta$.
@@
We first consider the case $i=0$. Let $v$ be in the image of $\varphi_{0}$, and let $\tilde{v}$ be a lift of $v$ in $\mathcal{L}(N(F),\Theta)^{\sigma}$. There exists a compact open subgroup $\mathcal{N}\subseteq N(F)$ such that
@@
Since $N(F)$ admits a filtration of $\Gamma$ stable compact open subgroups, we assume that both $\mathcal{N}$ and $\mathcal{N}^{\prime}$ are $\Gamma$-stable. If $X$ denotes the coset space $\mathcal{N}/\mathcal{N}^{\prime}$, then we have
@@
Since the action of $\Gamma$ on $X\setminus X^{\sigma}$ is free, there exists a subset $\mathcal{U}\subseteq X\setminus X^{\sigma}$ such that $X\setminus X^{\sigma}$ is the disjoint union of $\sigma^{i}\mathcal{U}$ for $1\leq i\leq\ell$. As $\tilde{v}$ is a $\sigma$-fixed vector, we get
@@
where ${\rm Nr}=1+\sigma+\cdots+\sigma^{\ell-1}\in\Lambda[\Gamma]$. This shows that
@@
This implies that the vector $v$ belongs to the Tate cohomology space $\widehat{H}^{0}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Conversely, let $w$ be an element of $\widehat{H}^{0}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Let $\mathcal{N}_{\sigma}$ be the compact open subgroup of $N^{\sigma}(F)$ with
@@
Now, consider the case $i=1$. Let $v$ be an element in the image of $\varphi_{1}$. Let $\tilde{v}$ be a lift of $v$ in ${\rm Ker}({\rm Nr}_{|\mathcal{L}(N(F),\Theta)})$. Then there exists $\Gamma$ stable compact open subgroups $\mathcal{N}^{\prime}\subset\mathcal{N}$ such that $\mathcal{N}^{\prime}$ fixes the vector $\tilde{v}$, the character $\Theta$ is trivial on $\mathcal{N}^{\prime}$, and
@@
where $X$ denotes the quotient space $\mathcal{N}/\mathcal{N}^{\prime}$. Since the $\Gamma$ action on $X\setminus X^{\sigma}$ is free, there exists a subset $\mathcal{V}\subseteq X\setminus X^{\sigma}$ such that $X\setminus X^{\sigma}=\coprod_{i=0}^{\ell-1}\sigma^{i}\mathcal{V}$. Then we have the identity
@@
Thus, we get that $\sum_{x\in X^{\sigma}}\overline{\Theta}^{-1}(x)\,\pi(x)v$ is trivial in $\widehat{H}^{1}(\mathcal{L})$, and this implies that $v\in\widehat{H}^{1}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Conversely, let $v$ be an element in $\widehat{H}^{1}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Let $\tilde{v}\in{\rm Ker}({\rm Nr})$ be a lift of $v$. Consider the element
@@
Theorem 3.3.2.
Let $P$ be a $\sigma$ stable $F$-parabolic subgroup of $G$, and let $N$ be the unipotent radical of $P$. Let $\mathcal{L}$ be a smooth $\Lambda[G(F)\rtimes\Gamma]$ module. Then the natural map $\widehat{H}^{i}(\mathcal{L})_{N^{\sigma}(F),\overline{\Theta}}\rightarrow\widehat{H}^{i}(\mathcal{L}_{N(F),\Theta})$ is injective.
@@
The long exact sequence of Tate cohomology corresponding to the above short exact sequence gives the following eact sequence:
@@
Using Lemma 3.3.1, the image of $\varphi_{i}$ is equal to $\widehat{H}^{i}(\mathcal{L})(N^{\sigma}(F),\overline{\Theta})$. Hence, we get the theorem. ∎
@@
Corollary 3.3.3.
Let $G$ be a connected reductive algebraic group defined over $F$, and let $\sigma$ be an order $\ell$ automorphism of $G$, where $\ell$ and $p$ are distinct primes. Let $\Pi$ be an $\ell$-modular cuspidal representation of $G(F)$ such that $\Pi$ extends to a representation of $G(F)\rtimes\Gamma$. Then each Tate cohomology group $\widehat{H}^{i}(\Pi)$ is a cuspidal representation of $G^{\sigma}(F)$.
@@
Let $Q$ be a proper $F$-parabolic subgroup of $G^{\sigma}$. By [KP23, Lemma 12.5.4], there exists a proper $\sigma$ stable $F$-parabolic subgroup $P$ of $G$ such that $P^{\sigma}=Q$. Let $N$ be the unipotent radical of $P$. Then $N^{\sigma}$ is equal to the unipotent radical of $Q$, to be denoted by $U$. Now, using Theorem 3.3.2 for $\Theta=1$, the trivial character of $N(F)$, we get an injection
@@
In the paper [TV16, Conjecture 6.1], the authors conjectured that if $\Pi$ is an irreducible $\ell$-modular representation of $G(F)$ such that $\Pi$ extends to a representation of $G(F)\rtimes\Gamma$, then the Tate cohomology groups $\widehat{H}^{i}(\Pi)$, $i\in\{0,1\}$, are finite length representations of $G^{\sigma}(F)$. In our previous work [ND24, Proposition 5.5], this conjecture is proved for $G={\rm GL}_{n}$ where $\sigma$ is induced by a Galois automorphism. In this subsection, we generalise this result for arbitrary connected reductive groups.
@@
Let $E/F$ be a finite Galois extension with $[E:F]=\ell$, where $\ell$ and $p$ are distinct primes. We fix a generator $\sigma$ of the cyclic group ${\rm Gal}(E/F)$. Let $G$ be a connected reductive group defined over $F$. Fix a minimal $F$-parabolic subgroup $P_{0}$ of $G$. Let $\Omega$ denote the set of proper $F$-parabolic subgroups of $G$ containing $P_{0}$. For $P\in\Omega$, we denote by $U$ the unipotent radical of $P$. Let $\pi$ be a smooth $R$-representation of $G(F)$, where the coefficient field $R$ is either $\overline{\mathbb{Q}}_{\ell}$ or $\overline{\mathbb{F}}_{\ell}$. We have the natural map obtained from Frobenius reciprocity:
$\pi\longrightarrow\prod_{P\in\Omega}i_{P(F)}^{G(F)}\circ r_{U(F)}(\pi),$ (3.3)
@@
$\pi_{c}\longrightarrow\prod_{P\in\Omega}i_{P(F)}^{G(F)}\circ r_{U(F)}(\pi_{c}).$ (3.4)
Since the functor $i_{P(F)}^{G(F)}$ preserves injective maps and $r_{U(F)}$ is an exact functor, we get a contradiction to the fact that $\pi_{c}$ is the kernel of the map (3.3). Thus, $\pi_{c}$ is a cuspidal representation. Using this observation, we now prove the following finiteness result.
Theorem 3.4.1.
Let $E/F$ be a finite Galois extension of prime degree $\ell$, where $\ell\neq p$. Let $G$ be a connected reductive group defined over $F$. Let $\Pi$ be a finite length $\ell$-modular representation of $[G(E)\rtimes{\rm Gal}(E/F)]$. Then, the Tate cohomology groups $\widehat{H}^{i}({\rm Gal}(E/F),\Pi)$ are finite length $\ell$-modular representations of $G(F)$.
@@
We use induction on the length of $\Pi$ and the dimension of $G$ to prove the theorem. All Tate cohomology groups are with respect to the action of the Galois group ${\rm Gal}(E/F)$, and we drop it from the notation of Tate cohomology. Suppose $\Pi$ is irreducible. Consider the natural map
$0\longrightarrow\widehat{H}^{i}(\Pi)_{c}\longrightarrow\widehat{H}^{i}(\Pi)\longrightarrow\prod_{P\in\Omega}i_{P(F)}^{G(F)}\circ r_{U(F)}(\widehat{H}^{i}(\Pi))$ (3.5)
For each parabolic $P\in\Omega$ with unipotent radical $U$ and an $F$-rational Levi subgroup $M$, the natural map $\Pi\rightarrow r_{U(E)}(\Pi)$ gives an $M(F)$ equivariant injective map (Theorem 3.3.2)
$r_{U(F)}(\widehat{H}^{i}(\Pi))\hookrightarrow\widehat{H}^{i}(r_{U(E)}(\Pi)).$
Since $r_{U(E)}(\Pi)$ is a finite length representation of $M(E)$ and ${\rm dim}(M)<{\rm dim}(G)$, the Tate cohomology $\widehat{H}^{i}(r_{U(E)}(\Pi))$ has finite length as a representation of $M(F)$ (by induction hypothesis). This implies that $r_{U(F)}(\widehat{H}^{i}(\Pi))$ is a finite length $M(F)$ representation, and therefore $i_{P(F)}^{G(F)}(r_{U(F)}(\widehat{H}^{i}(\Pi)))$ is a finite length $G(F)$ representation as the parabolic induction preserves finite length ([Vig96, Chapter II, Section 5.13]). Thus, the representation
$\prod_{P\in\Omega}i_{P(F)}^{G(F)}\circ r_{U(F)}(\widehat{H}^{i}(\Pi))$
is of finite length. Note that the center $Z(E)$ of $G(E)$ acts on the space of $\Pi$ by a character (as $\Pi$ is irreducible). We deduce that the center $Z(F)$ of $G(F)$ acts on $\widehat{H}^{i}(\Pi)$ by a character. Hence, the cuspidal $G(F)$ representation $\widehat{H}^{i}(\Pi)_{c}$, being a subrepresentation of $\widehat{H}^{i}(\Pi)$, admits a central character. Since $\widehat{H}^{i}(\Pi)$ is finitely generated as a representation of $G(F)$ (Theorem 2.3.3), it follows from [Dat09, Corollary 4.5]) that $\widehat{H}^{i}(\Pi)_{c}$ is also finitely generated. Thus, $\widehat{H}^{i}(\Pi)_{c}$ is a finitely generated $\ell$-modular cuspidal representation of $G(F)$ with a central character and we get that $\widehat{H}^{i}(\Pi)_{c}$ is of finite length ([Vig96, Section 7.4, Chapter I]). We conclude from (3.5) that the $\ell$-modular $G(F)$ representation $\widehat{H}^{i}(\Pi)$ is of finite length.
Now, consider the general case where $\Pi$ is not necessarily an irreducible representation of $\overline{\mathbb{F}}_{\ell}[G(F)]$. Let $\sigma$ be a generator of $\Gamma={\rm Gal}(E/F)$. Let $\Pi_{1}$ be an irreducible sub-representation of $\Pi$ and set $\Pi_{1}^{\prime}$ to be the sub-representation of $\Pi$ spanned by $\Pi_{1}^{\sigma^{i}}$, for $0\leq i<\ell$. Note that $\Pi_{1}^{\prime}$ is $\sigma$-stable, and we consider the short exact sequence of $\overline{\mathbb{F}}_{\ell}[G(E)\rtimes\Gamma]$ modules
@@
As a corollary, we get the following result.
Corollary 3.4.2.
Let $E/F$ be a finite Galois extension of prime degree $\ell$, where $\ell\neq p$. Let $G$ be a connected reductive group defined over $F$. Let $\Pi$ be a finite length integral $\mathcal{K}$ representation of $G(E)\rtimes{\rm Gal}(E/F)$. Then, for any $\Lambda[G(E)\rtimes{\rm Gal}(E/F)]$ stable lattice $\mathcal{L}$ in $\Pi$, the Tate cohomology spaces $\widehat{H}^{i}({\rm Gal}(E/F),\mathcal{L})$ are finite length $\ell$-modular representations of $G(F)$.
Proof.
Recall that $\mathcal{L}$ is a free $\Lambda$-module and that $\mathcal{L}/\ell\mathcal{L}$ is a finite length representation of $G(E)$ (see [Vig96, II.5.11.a] for finiteness). Consider the short exact sequence of $G(E)\rtimes{\rm Gal}(E/F)$ modules
$0\longrightarrow\mathcal{L}\xrightarrow{{\rm mult.}\ell}\mathcal{L}\longrightarrow\mathcal{L}/\ell\mathcal{L}\longrightarrow 0$
By Theorem 3.4.1, the Tate cohomology groups $\widehat{H}^{i}(\mathcal{L}/\ell\mathcal{L})$ are of finite length as representations of $G(F)$. From the long exact sequence of Tate cohomology corresponding to the above short exact sequence, we have
$0\rightarrow\widehat{H}^{0}(\mathcal{L})\rightarrow\widehat{H}^{0}(\mathcal{L}/\ell\mathcal{L})\rightarrow\widehat{H}^{1}(\mathcal{L})\rightarrow 0$
and
$0\rightarrow\widehat{H}^{1}(\mathcal{L})\rightarrow\widehat{H}^{1}(\mathcal{L}/\ell\mathcal{L})\rightarrow\widehat{H}^{0}(\mathcal{L})\rightarrow 0$
Thus, we get that each $\widehat{H}^{i}(\mathcal{L})$ is of finite length. ∎
4. Some genericity and irreducibility results
In this section, $G$ will be a split connected reductive algebraic group defined over $F$. We follow the notations defined in Subsection 3.4.1.
4.0.1.
For this section, we follow [BH02]. Let $B$ be a Borel subgroup of $G$ defined over $F$. Let $E$ be a cyclic Galois extension of $F$ with $[E:F]=\ell$. Let $S$ be a maximal $F$-split torus and let $\Phi$ the set of roots of $S$ in $G$, and let $\Delta$ be the set of simple roots in $\Phi$ for the choice of $B$. Let $\mathfrak{g}$ be the Lie algebra of $G$ and let $\mathfrak{g}_{\alpha}$ be the subspace of $\mathfrak{g}$ for which $S$ acts via $\alpha$. The root subgroup corresponding to $\alpha\in\Delta$ is denoted by $U_{\alpha}$, this is the unique connected subgroup of $G$ whose Lie algebra is $\mathfrak{g}_{\alpha}$. Let $R$ denote $\overline{\mathbb{Q}}_{\ell}$ or $\overline{\mathbb{F}}_{\ell}$. A character $\Theta:U(F)\rightarrow R^{\ast}$ is called non-degenerate whenever $\Theta$ is non-trivial on $U_{\alpha}(F)$ for all $\alpha\in\Delta$. Note that
$U(E)/U(E)_{\rm der}\simeq\prod_{\alpha\in\Delta}V_{\alpha}(E),$
where $V_{\alpha}$ is the one-dimensional $F$-vector group. Thus we choose a non-degenerate character $\Theta_{E}:U(E)\rightarrow R^{\ast}$, which is ${\rm Gal}(E/F)$ stable and the restriction of $\Theta_{E}$ to $U(F)$ is denoted by $\Theta_{F}$ (which is a non-degenerate character of $U(F)$). We set